# DFREC: DeepFake Identity Recovery Based on Identity-aware Masked Autoencoder

### v1

@@ IV-A3 Evaluation Metrics
We utilize the Frechet Inception Distance (FID) and Identity Similarity (IDSim) metrics to assess the identity recovery performance. FID is commonly employed for evaluating image similarity, while IDSim is specifically designed to assess the identity similarity between the reconstructed image and the original image [39]. To compute IDSim, we leverage the pretrained FaceNet model to extract identity features from both the original and recovered target faces, followed by computing the cosine similarity between these extracted features. Additionally, we use $Acc_{id}$ to represent the identity-based recovery accuracy. It is computed based on the identity similarity between the input face and the recovered target face. A high identity similarity implies a small discrepancy, and vice versa. If the identity similarity surpasses the threshold of 0.9, the input face is categorized as real. Otherwise, the input face is classified as fake.
@@ IV-A4 Implementation Details
IV-B Deepfake Recovery Results Analysis
@@
Quantitative Evaluation We use FID to assess the image recovery quality, and IDSim to evaluate identity recovery quality between the recovered and the original target face for comparison with other methods. We leverage publicly available pre-trained models or existing codes for MAT, Repaint, and RECCE. Delocate and DFI are implemented by following the detailed methodologies presented in the corresponding papers. For face inpainting methods like MAT and Repaint, we employ dlib for the facial detection, and then mask areas that contain facial components before inpainting. Table I presents the recovery performance across the three datasets. The ‘None’ method in the first line shows the FID scores between the forgery face and the original target face, and the next six lines show the values of the corresponding metrics between the recovered target face and the original target face. We analyze the results of image recovery on the FaceForensics++, CelebaMegaFS, and FFHQ-E4S datasets. Compared with existing face inpainting methods [37, 38] and deepfake recovery methods [23, 24, 26], DFREC achieves the best FID scores and identity similarity in different types of face swapping forgeries, demonstrating a higher capability of generating images that are closer to the original target face identity.
@@
Qualitative Evaluation Figures 3 - 5 compare the target identity recovery quality between DFREC and several SOTA face inpainting and deepfake recovery methods. The evaluations were conducted on the test sets of the FaceForensics++, CelebaMegaFS, and FFHQ-E4S, respectively. As shown in Figures 3 - 5, although MAT and Repaint can synthesize realistic facial images based on the unmasked region, there are still identity gaps between the synthesized images and the original target faces. We also compared DFREC with three deepfake recovery approaches [23, 24, 26]. They tend to generate blurred and unidentified faces, which are apparently different from the original target faces. In contrast, DFREC can synthesize more realistic target faces with similar identities as the original target faces.
IV-C Generalization Evaluation of Deepfake Recovery
@@
In this section, we assess the generalized deepfake recovery performance of DFREC on the FaceForensics++ and CelebaMegaFS datasets. We trained DFREC on the CelebaMegaFS dataset and then evaluate its deepfake recovery accuracy across the two datasets. Table II presents the identity-based recovery accuracy $Acc_{id}$ on various types of forgery data. It can be observed that DFREC generally outperforms DFI [26] in terms of the generalized deepfake recovery. To visually demonstrate the generalization effect of DFREC, we present the recovered target images across different datasets in Figure 6. It can be seen that DFREC could well recover the target faces from the unseen forgery images.
@@
TABLE II: Generalization of deepfake recovery compared with DFI [26] across diverse forgery images. $Acc_{id}$ is used as the evaluation metric. Intra-dataset evaluations are highlighted in red.
@@
Figure 6: Target face recovery on unseen forgeries. The vertical axis represents the training dataset, and the horizontal axis denotes the evaluation dataset. From top to bottom on the left side of the recovered images are the forged face, target face, source face, and segmentation map. The value at the top-right corner indicates the identity similarity between the recovered and target faces.
@@
IV-D Ablation Study
@@
Identity Segmentation Module To demonstrate its significance, we remove the ISM and input the image directly to the TIRM. For the MAE of TIRM, we employ a uniform mask encompassing the entire target face region. The visualization results on the CelebaMegaFS dataset are shown in Figure 7. It shows that DFREC with ISM tends to generate more accurate target face. We also conducted quantitative evaluations on the three types of forgeries with CelebaMegaFS. The identity similarity between the recovered and the original target face is calculated and shown in Table III. The effectiveness of the ISM is corroborated by the improved identity similarity when the ISM is included.
Target Identity Fusion In DFREC, we hypothesize that the segmented source information contains target identity features. Thus, we use the SIRM to extract the target identity features and fuse them with the visible token embeddings obtained from the MAE to facilitate the target face recovery. To demonstrate its efficacy, we conduct an ablation study by removing the fusion strategy to solely rely on unmasked patches for recovery. As shown in Table III, the ability of DFREC to recover the target face is relatively weak without the fusion strategy. The improvement is significant with the fusion strategy, which attests to our hypothesis that the identity fusion plays a pivotal role in enhancing the quality of the target face recovery.
TABLE III: Ablation experiments of the ISM and Fusion strategy. The evaluations are conducted on the CelebaMegaFS images deepfaked by FTM, IDInjection and LCR.
@@
V Conclusion

### v2

@@ IV-A3 Evaluation Metrics
We utilize the Frechet Inception Distance (FID) and Identity Similarity (IDSim) metrics to assess the identity recovery performance. FID is commonly employed for evaluating image similarity, while IDSim is specifically designed to assess the identity similarity between the reconstructed image and the original image [40]. To compute IDSim, we leverage the pretrained FaceNet [30] model to extract identity features from both the original and recovered target faces, followed by computing the cosine similarity between these extracted features. Attribute preservation is measured by computing cosine similarity between adaptive attribute embeddings [34]. Pose consistency is evaluated using pose parameters from a 3D face reconstruction model [41] with angular distance measurement. Additionally, we use $Acc_{id}$ to represent the identity-based recovery accuracy. It is computed based on the identity similarity between the input face and the recovered target face. A high identity similarity implies a small discrepancy, and vice versa. If the identity similarity surpasses the threshold of 0.9, the input face is categorized as real. Otherwise, the input face is classified as fake.
@@ IV-A4 Implementation Details
IV-B Intra-dataset Deepfake Recovery
@@
IV-B1 Quantitative Evaluation
We use FID to assess the image recovery quality, and IDSim to evaluate identity recovery quality between the recovered and the original target face for comparison with other methods. We leverage publicly available pre-trained models or existing codes for MAT, Repaint, and RECCE. Delocate and DFI are implemented by following the detailed methodologies presented in the corresponding papers. For face inpainting methods like MAT and Repaint, we employ dlib for the facial detection, and then mask areas that contain facial components before inpainting. Table I presents the recovery performance across the three datasets. The ‘None’ method in the first line shows the FID and IDSim scores between the forgery face and the original target face, and the next six lines show the values of the corresponding metrics between the recovered target face and the original target face. We analyze the results of image recovery on the FaceForensics++, CelebaMegaFS, and FFHQ-E4S datasets. Compared with existing face inpainting methods [38, 39] and deepfake recovery methods [24, 25, 27], DFREC achieves the best FID scores and identity similarity in different types of face swapping forgeries, demonstrating a higher capability of recovering images that are closer to the original target face identity.
@@ IV-B2 Qualitative Evaluation
Figs 3 - 5 compare the target identity recovery quality between DFREC and several SOTA face inpainting and deepfake recovery methods. The evaluations were conducted on the test sets of the FaceForensics++, CelebaMegaFS, and FFHQ-E4S, respectively. As shown in Figs 3 - 5, although MAT and Repaint can synthesize realistic facial images based on the unmasked region, there are still identity gaps between the synthesized images and the original target faces. We also compared DFREC with three deepfake recovery approaches [24, 25, 27]. They tend to generate blurred and unidentified faces, which are apparently different from the original target faces. In contrast, DFREC can synthesize more realistic target faces with similar identities as the original target faces.
IV-C Generalization of Deepfake Recovery
@@
In this section, we evaluate the deepfake recovery performance of DFREC on unseen forgeries. To visually demonstrate the generalization effect of DFREC, we present the recovered target images across different datasets in Fig. 6. DFREC are trained on the various forgery types of CelebaMegaFS dataset and then evaluated across FaceForensics++ and CelebaMegaFS datasets. It can be seen that DFREC could well recover the target faces from the unseen forgery images. Also, we calculated the identity between the recovered target faces and the original target faces. DFREC is trained on a specific forgery set and then evaluated across other forgeries. Table II presents the identity similarity IDSim on various types of forgery data. It can be observed that DFREC generally outperforms DFI [27] in terms of the generalized deepfake recovery.
@@
TABLE II: Generalization of deepfake recovery compared with DFI [27] across diverse forgery images in the CelebaMegaFS dataset. $IDSim$ is used as the evaluation metric. Intra-dataset evaluations are highlighted in red.
Methods Train FTM IDI LCR
DFI [27] 0.4137 0.4318 0.3098
DFREC FTM 0.7192 0.6586 0.4711
DFI [27] 0.3938 0.4346 0.2966
DFREC IDI 0.6203 0.7582 0.4546
DFI [27] 0.3603 0.3802 0.3442
DFREC LCR 0.6332 0.6843 0.6967
Figure 6: Target face recovery on unseen forgeries. The vertical axis represents the training dataset, and the horizontal axis denotes the evaluation dataset. From top to bottom of the inset on the left side of the recovered images are the forged face, target face, source face, and segmentation map. The value at the top-right corner indicates the identity similarity between the recovered and target faces.
IV-D Deepfake Recovery Accuracy
The proposed DFREC should exhibit distinct behaviors when processing manipulated versus authentic images. As illustrated in Fig. 7, DFREC generates different output results when presented with forgery and pristine facial images. Specifically, for forgery faces, DFREC precisely reconstructs both the source and target faces from the manipulated input. In contrast, for pristine faces, DFREC ensures that the recovered source and target faces remain identical to the original input. To further evaluate the effectiveness of our approach, we calculated the recovery accuracy ($Acc_{id}$) under intra-dataset and cross-dataset settings.
Figure 7: Intermediate outputs of DFREC alongside the reconstructed source and target faces. The segmentation map $m_{s}$ is generated by the ISM, and the masked image is produced by applying $m_{s}$ to the input image. Our DFREC could effectively distinguishes between forgery and pristine regions, enabling precise recovery of both source and target faces.
IV-D1 Intra-dataset Recovery Accuracy
We first assess the intra-dataset recovery accuracy of DFREC on the FaceForensics++, CelebaMegaFS, and FFHQ-E4S datasets. The identity similarity between the recovered target image and the forged image should exhibit significant differences, whereas for authentic images, the similarity should remain consistent. Table III presents the identity-based detection accuracy $Acc_{id}$ on various types of forgery and pristine images. We compare DFREC with existing deepfake recovery methods [24, 25, 27] across three datasets. It can be observed that DFREC generally outperforms existing deepfake recovery methods in terms of detection performance. Its face recovery accuracy is obviously higher than RECCE[24], Delocate[25], and DFI[27].
TABLE III: Comparison of $Acc_{id}$ with deepfake recovery methods on the FaceForensics++, CelebaMegaFS, and FFHQ-E4S datasets.
Methods FaceForensic++ CelebaMegaFS FFHQ-E4S
DF FShi LCR FTM IDI LCR E4S
RECCE [24] 43.39 49.70 48.77 52.25 56.41 56.00 50.35
Delocate [25] 89.32 89.83 67.16 67.37 68.82 65.52 69.01
DFI [27] 90.81 86.50 94.43 96.45 92.70 98.83 97.18
DFREC 100 99.38 99.55 99.13 99.32 99.07 98.04
IV-D2 Cross-dataset Recovery Accuracy
Here, we assess the generalized deepfake recovery performance of DFREC. We trained DFREC on the CelebaMegaFS dataset and then evaluate its deepfake recovery accuracy across FaceForensics++ and CelebaMegaFS datasets. Given that DFI is specifically designed for deepfake recovery and represents the best-performing existing method, we compare the deepfake recovery performance of DFREC against DFI. Table IV presents the identity-based recovery accuracy $Acc_{id}$ on various types of forgery data. It can be observed that DFREC generally outperforms DFI [27] in terms of the generalized deepfake recovery.
TABLE IV: Generalization of deepfake recovery compared with DFI [27] across diverse forgery images. $Acc_{id}$ is used as the evaluation metric. Intra-dataset evaluations are highlighted in red.
@@ IV-D2 Cross-dataset Recovery Accuracy
IV-E Robustness Evaluation of Deepfake Recovery
TABLE V: Robustness evaluation of three forgery types (FTM, IDInjection, and LCR) in the CelebaMegaFS dataset.
Methods GaussianBlur JPEGCompression GaussianNoise
FTM IDI LCR FTM IDI LCR FTM IDI LCR
DFI [27] 0.3788 0.4230 0.3294 0.3999 0.4394 0.3421 0.3864 0.4244 0.3366
DFREC 0.6902 0.7240 0.6761 0.7055 0.7305 0.6816 0.7039 0.7324 0.6793
We also evaluated the robustness of our approach on the CelebaMegaFS dataset. Specifically, we applied various post-processing operations to the input forgery images and subsequently measured the identity similarity between the recovered target images and the original target images, as shown in Table V. In this context, GB refers to Gaussian Blur (with a random kernel size ranging from 3 to 7), JPEG indicates JPEG Compression (with a random quality setting between 50 and 99), and GN represents Gaussian Noise (with variance randomly set between 10 and 50). Compared with DFI [27], our method exhibits better robustness to these perturbations, demonstrating its suitability for real-world applications.
IV-F Ablation Study
@@
IV-F1 Identity Segmentation Module
To demonstrate its significance, we remove the ISM and input the image directly to the TIRM. For the MAE of TIRM, we employ a uniform mask encompassing the entire target face region. The visualization results on the CelebaMegaFS dataset are shown in Fig. 8. It shows that DFREC with ISM tends to generate more accurate target face. We also conducted quantitative evaluations on the three types of forgeries with CelebaMegaFS. The identity similarity between the recovered and the original target faces is calculated and shown in Table VI. The effectiveness of the ISM is corroborated by the improved identity similarity when the ISM is included.
IV-F2 Target Identity Fusion
In DFREC, we hypothesize that the segmented source information contains target identity features. Thus, we use the SIRM to extract the target identity features and fuse them with the visible token embeddings obtained from the MAE to facilitate the target face recovery. To demonstrate its efficacy, we conduct an ablation study by removing the fusion strategy to solely rely on unmasked patches for recovery. As shown in Table VI, the ability of DFREC to recover the target face is relatively weak without the fusion strategy. The improvement is significant with the fusion strategy, which attests our hypothesis that the identity fusion plays a pivotal role in enhancing the quality of the target face recovery.
TABLE VI: Ablation experiments of the ISM and Fusion strategy. The evaluations are conducted on the CelebaMegaFS images deepfaked by FTM, IDInjection and LCR.
@@ IV-F2 Target Identity Fusion
IV-F3 Effect of Each Loss
In our approach, we incorporate identity loss ($\mathcal{L}_{id}$), perceptual loss ($\mathcal{L}_{perc}$), patch loss ($\mathcal{L}_{patch}$), and attribute loss ($\mathcal{L}_{attr}$) to ensure the quality of both the reconstructed source and target faces. To evaluate the effectiveness of each loss component, we trained and tested the model on the FTM subset of the CelebaMegaFS dataset using various loss configurations. The experimental results (Identity Similarity, Attribute Consistency, Pose Preservation, PSNR, and SSIM) are summarized in Table VII. All four loss terms contribute to performance gains. We take the model using only identity loss as our baseline, which achieves satisfactory identity preservation but suffers from lower visual quality. The patch loss then refines pixel-level details, and adding perceptual loss further enhances visual metrics (PSNR, SSIM). Finally, incorporating the attribute loss improves the accuracy of facial attributes and poses (Attr, Pose). By jointly leveraging these four loss functions, our method achieves superior reconstruction of the target face, striking a favorable balance between identity fidelity and visual realism.
TABLE VII: Ablation experiments of each loss. The evaluations are conducted on the CelebaMegaFS images deepfaked by FTM, IDInjection and LCR.
$\mathcal{L}_{id}$ $\mathcal{L}_{perc}$ $\mathcal{L}_{patch}$ $\mathcal{L}_{attr}$ IDSIM Attr Pose PSNR SSIM
$\checkmark$ 0.6926 90.14 130.65 12.23 0.4239
$\checkmark$ $\checkmark$ 0.7004 90.70 108.98 12.27 0.4262
$\checkmark$ $\checkmark$ 0.7072 94.96 10.34 21.93 0.6701
$\checkmark$ $\checkmark$ 0.6390 95.54 4.89 21.59 0.6946
$\checkmark$ $\checkmark$ $\checkmark$ 0.6336 95.25 4.27 21.88 0.6936
$\checkmark$ $\checkmark$ $\checkmark$ 0.6427 95.28 5.10 21.40 0.6715
$\checkmark$ $\checkmark$ $\checkmark$ 0.7109 95.02 10.09 21.92 0.6758
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.7192 96.07 2.62 22.50 0.7507
V Limitations
In this work, we introduce a deepfake face recovery method based on masked autoencoders, capable of reconstructing both the original target face and the source face from a face-swapped image. Despite its effectiveness, our approach has some limitations. First, deepfake face recovery is inherently challenging, and the success of our method relies on the presence of identity features related to the target image within the forged image. In scenarios where the forged face is not derived from the target face (e.g., through direct cropping and splicing), the performance of our approach is significantly constrained. Second, substantial variations in face images due to differing capture conditions can affect the performance of our method. This limitation can be mitigated by assembling a sufficiently diverse set of face images and employing existing deepfake face-swapping algorithms to generate synthetic data for training. Nevertheless, our proposed method can effectively recover both the original source and target faces, offering a novel tool for assisting in the collection of judicial evidence.
VI Conclusion
@@
Biography Section
[Uncaptioned image] Peipeng Yu is currently pursuing Ph.D. degree in College of Information Science and Technology/Cyber Security at Jinan University, Guang Zhou, China. His research interests include artificial intelligence security and video forensics.
[Uncaptioned image] Hui Gao is currently pursuing master’s degree in College of Cyber Security at Jinan University, Guang Zhou, China. His research interests include deepfake detection and image forensics.
[Uncaptioned image] Zhitao Huang is currently pursuing master’s degree in College of Information Science and Technology/Cyber Security at Jinan University, Guang Zhou, China. His research interests include artificial intelligence security and backdoor attacks.
[Uncaptioned image] Zhihua Xia received his Ph.D. degree in computer science and technology from Hunan University, China, in 2011, and worked successively as a lecturer, an associate professor, and a professor with College of Computer and Software, Nanjing University of Information Science and Technology. He is currently a professor with the College of Cyber Security, Jinan University, China. He was a visiting scholar at New Jersey Institute of Technology, USA, in 2015, and was a visiting professor at Sungkyunkwan University, Korea, in 2016. He serves as a managing editor for IJAACS. His research interests include AI security, cloud computing security, and digital forensic. He is a member of the IEEE since Mar. 1, 2014.
[Uncaptioned image] Chip-Hong Chang (S’92-M’98-SM’03-F’18) received the B.Eng. (Hons.) degree from the National University of Singapore in 1989, and the M. Eng. and Ph.D. degrees from Nanyang Technological University (NTU) of Singapore in 1993 and 1998, respectively. He is a Professor of the School of Electrical and Electronic Engineering (EEE) of NTU. He held joint appointments with the university as Assistant Chair of Alumni from 2008 to 2014, Deputy Director of the Center for High Performance Embedded Systems from 2000 to 2011, and Program Director of the Center for Integrated Circuits and Systems from 2003 to 2009. He has coedited 6 books, published 13 book chapters, more than 120 international journal papers (>90 are in IEEE) and more than 200 refereed international conference papers (mostly in IEEE), and delivered over 60 keynotes, tutorials and invited seminars. His current research interests include hardware security, AI security, biometric security, trustworthy sensing and hardware accelerators for post-quantum cryptography and edge computational intelligence. Dr. Chang currently serves as the Senior Area Editor of IEEE Transactions on Information Forensic and Security (TIFS) and IEEE Journal on Emerging and Selected Topics in Circuits and Systems, and Associate Editor of IEEE Transactions on Very Large Scale Integration (VLSI) Systems. He also served as the Associate Editor of TIFS, IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems, IEEE Access, and IEEE Transactions on Circuits and Systems-I. He is an IEEE Fellow, IET Fellow, AAIA Fellow and the 2018-2019 IEEE Circuits and Systems Society Distinguished Lecturer.